Grade 9Mechanics

a bolck lying ona inclined plane has a weight os 50N .It is just begins to slide down when the inclination of plane with the horizontal is 30 The value of U s is

uday surya
6 Years agoGrade 9

4 Answers

Arun
6 Years ago
Dear Uday
 
I didn’t get U. Please elaborate it or you can also attach an image. I will be happy to help you.
Khimraj
6 Years ago

The block reaches the point of sliding when the plane makes an angle of 30º with the horizontal.

Hence in this situation, frictional force = µsN

Balancing the forces :

N = W cos 30º

µsN = W sin 30º

µs W cos 30º = W sin 30º

µs = 1/\sqrt{3}

Khimraj
6 Years ago

The block reaches the point of sliding when the plane makes an angle of 30º with the horizontal.

Hence in this situation, frictional force = µsN

Balancing the forces :

N = W cos 30º

µsN = W sin 30º

µs W cos 30º = W sin 30º

µs = 1/\sqrt{3}

aswanth nayak
6 Years ago

Dear Student,

The block reaches the point of sliding when the plane makes an angle of 30º with the horizontal.

Hence in this situation, frictional force = µsN

Balancing the forces :

N = W cos 30º

µsN = W sin 30º

µs W cos 30º = W sin 30º

µs = 1/\sqrt{3}

 

Hope it helps you

Regards